OpenStudy (anonymous):

-x+y=1 x+y=0 who know how to solve this problem and graph it i need helpppppppppppppppppp

OpenStudy (jdoe0001):

solve for "x" on each equation then graph them both

OpenStudy (zzr0ck3r):

I think they want you to solve then graph you have 2 equations x+y = 0 and -x+y = 1 solving for y on the left y = -x solving for y on the right y = 1+x set them equal to each other and solve for y